You are offered $100,000 today or $313,000 in 14 years. Assuming that you can earn 1 1% on your money, which should you choose? Click the table icon to view the PVIF table EEE The present value of the $313,000 in 14 years is S (Round to the nearest cent)
Answer #1

Present Value of $313,000 in 14 Years

Present Value = Future Value x Present Value Factor

= $313,000 x [PVIF 11%, 14 Years]

= $313,000 x [1/(1 + r)n]

= $313,000 x [1/(1 + 0.11)14]

= $313,000 x [1/4.31044]

= $313,000 x 0.23199

= $72,614.38

“Present Value of $313,000 in 14 Years = $72,614.38”

DECISION

We should choose the receipt of $100,000 today, since the Present Value of $313,000 in 14 years ($72,614.38) is less than the amount to be received today ($100,000)
Answer #2

SOLUTION :


It requires to compare the present values of options to select the better one. Higher the PV, better is the option. 


Accordingly, we determine PVs of the two options .


Option 1 :


$100000 today .

Its PV = $100000 .


Option 2 :


Receiving $313000 in 14 years from now. 

Required rate, r = 11% = 0.11

=> 1 + r = 1.11


PV of Option 2 

= FV/(1 + r)^n 

= 313000 / 1.11^14 

= 72614.38 ($) 


PV of Option 2 is = $72614.38 (ANSWER)


So, PV of Option 1 is higher than that of Option 2. 

Hence, Option 1 should be accepted. (ANSWER).



answered by: Tulsiram Garg